OpenStudy (anonymous):

Simplify the expression: x-4/x^2-49 - 1/x+7 all over x/x-7

OpenStudy (anonymous):

hard to parse this one. maybe \[\frac{\frac{x-4}{x^2-49}-\frac{1}{x+7}}{\frac{x}{x-7}}\]\]

OpenStudy (anonymous):

thank you, yes that is how it should look.

OpenStudy (anonymous):

ok then if i were you i would multply top and bottom of the big fraction by \[\frac{x^2-49}{x^2-49}\] to clear the fractions

OpenStudy (anonymous):

with the understanding that \[x^2-49=(x+7)(x-7)\] so all the fractions will cancel and you will just get a numerator and denominator to combine like terms

OpenStudy (anonymous):

i will write the steps if you like

OpenStudy (anonymous):

that would help a lot, thanks. I'm kind of slow when it comes to this heh

OpenStudy (anonymous):

ok so lets go real slow.

OpenStudy (anonymous):

the denominator is \[\frac{x}{x-7}\] so if you multiply it by \[(x+7)(x-7)\] the \[x-7\] will cancel. you will get \[\frac{x}{x-7}\times (x+7)(x-7)=x(x+7)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so that will be our denominator. we are not done of course

OpenStudy (anonymous):

now multiply the numerator by \[(x+7)(x-7)=x^2-49\]

OpenStudy (anonymous):

you get \[(\frac{x-4}{x^2-49}-\frac{1}{x+7})\times (x+7)(x-7)\]

OpenStudy (anonymous):

for the first term everything cancels. for the second term the \[x+7\] cancels giving \[x-4-1(x-7)\]

OpenStudy (anonymous):

the parentheses are important because it is - the whole thing.

OpenStudy (anonymous):

so now we have a fraction as \[\frac{x-4-1(x-7)}{x(x+7)}\]

OpenStudy (anonymous):

alright, I'm still with you :3

OpenStudy (anonymous):

well we are essentially done. now it is just a matter of multiplying out and combining like terms. i get \[ \frac{x-4-x+7}{x^2+7x}\] or \[\frac{3}{x^2+7x}\]

OpenStudy (anonymous):

yay =) I got that too~ Also, do you mind helping me understand one more? it's a bit easier I think

OpenStudy (anonymous):

k

OpenStudy (anonymous):

it's x/x+y-2xy/x^2-y^2+y/x-y

OpenStudy (anonymous):

yikes

OpenStudy (anonymous):

sorry, I don't know how to write it pretty like you.

OpenStudy (anonymous):

\[\frac{x}{x+y}-\frac{2xy}{x^2-y^2}+\frac{y}{x-y}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

use \[x^2-y^2=(x+y)(x-y)\] as your common denominator.

OpenStudy (anonymous):

you have to multiply the first one top and bottom by \[x-y\] the middle one is fine as it is and the last one multiply top and bottom by \[x+y\] that should give it to you

OpenStudy (anonymous):

denominator will just be \[x^2-y^2\] numerator will be \[x(x-y)-2xy+y(x+y)\]

OpenStudy (anonymous):

multiply out and combine like terms. i get \[x^2-xy-2xy+xy+y^2\] \[x^2-2xy+y^2\] \[(x-y)(x-y)\] and i guess now you can cancel a \[x-y\] from top and bottom to get \[\frac{x-y}{x+y}\]

OpenStudy (anonymous):

awesome! That made much more sense <3 I was trying to figure out where the 2 went lol. Thank you so much for your help, is there a way to give you super bonus points?